jigglypuff314 (jigglypuff314):

A rectangular garden of area 75 ft^2 is bounded on 3 sides by a wall costing $8 per ft, and on the fourth by a fence costing $4 per ft. What are the most economical dimensions of the garden?

jigglypuff314 (jigglypuff314):

Could someone at least check this for me? pwease? A = L * W = 75 so L = 75/W Cost = 4W + 8(2L + W) = 4W + 16L + 8W = 12W + 16L = 12W + 1200/W C' = (12(W^2 - 100))/(W^2) = (12(W-10)(W+10))/(W^2) so W = 10 then L = 7.5 ?

jigglypuff314 (jigglypuff314):

@annas can you help? :3

OpenStudy (anonymous):

Just a min :)

OpenStudy (anonymous):

W = 10 is correct wait lemme check other

OpenStudy (anonymous):

oh yes L = 7.5 correct ^_^

OpenStudy (anonymous):

sorry I was also doing some other work xD

jigglypuff314 (jigglypuff314):

(fist pumpt) w00t! yes! :D Thanks :)